A particle is traveling along a straight path described as = 0.75x. If its position along the x-axis is x=(8t)m, where t is in seconds, determine its speed when t=2s.
1
Expert's answer
2020-08-10T19:58:35-0400


If

"x=8t (m),"

then

"v_x=dx\/dt=8 (m\/s)."

From the picture:

"v=v_x\/cos\u2061\u03b1 \n."

If the particle is travelling along a straight path "y=0.75x" , then "tan\u2061\u03b1=0.75." In this case . "\u03b1=atan\u20610.75\u224836.87\u00b0."


"v=8\/cos\u2061(36.87\u00b0) \u22488\/0.8=10 (m\/s)\n."

Answer: "v=10 (m\/s)"